# Graph Adapter of EEG Foundation Models for Parameter Efficient Fine Tuning

### v1

@@ Abstract
In diagnosing mental diseases from electroencephalography (EEG) data, neural network models such as Transformers have been employed to capture temporal dynamics. Additionally, it is crucial to learn the spatial relationships between EEG sensors, for which Graph Neural Networks (GNNs) are commonly used. However, fine-tuning large-scale complex neural network models simultaneously to capture both temporal and spatial features increases computational costs due to the more significant number of trainable parameters. It causes the limited availability of EEG datasets for downstream tasks, making it challenging to fine-tune large models effectively. We propose EEG-GraphAdapter (EGA), a parameter-efficient fine-tuning (PEFT) approach to address these challenges. EGA is integrated into pre-trained temporal backbone models as a GNN-based module and fine-tuned itself alone while keeping the backbone model parameters frozen. This enables the acquisition of spatial representations of EEG signals for downstream tasks, significantly reducing computational overhead and data requirements. Experimental evaluations on healthcare-related downstream tasks of Major Depressive Disorder and Abnormality Detection demonstrate that our EGA improves performance by up to 16.1% in the F1-score compared with the backbone BENDR model.
Introduction
@@
Electroencephalography (EEG) is a widely used technique for detecting brain activity, capturing electrical signals from various brain regions through electrodes. While other imaging methods, such as functional MRI (fMRI), offer complementary insights into brain function, EEG uniquely provides direct and real-time access to brain dynamics. These signals reveal information related to functions such as consciousness, cognition, and motor activity, with applications spanning healthcare fields, including diagnosing mental disorders and brain-computer interfaces (BCIs). As such, developing a foundational model for analyzing EEG data across diverse downstream tasks is essential in this domain.
@@
In recent years, with the development of large-scale neural network models such as LLMs, many foundation models processing time series data have been proposed, including Transformer-based models. Methods for applying these foundation models to EEG signals have also been proposed (Kostas, Aroca-Ouellette, and Rudzicz 2021; Wang et al. 2023a).
@@
However, there are several technical challenges in learning EEG representations using fusion models that combine multiple components: (1) Unaffordable computational cost: As the number of parameters for fine-tuning in each downstream task increases, the computational resources for downstream tasks become more significant. (2) Insufficient training data: The available labeled EEG data sets are limited for many downstream tasks particularly in the healthcare domain due to the considerable effort required for data acquisition, measurement, and labeling by doctors. If fine-tuning is performed with insufficient datasets, problems such as overfitting may arise, which prevents the required accuracy from being achieved in the downstream task. In particular, the demand for predictions is increasing in healthcare-related domains that predict neurological disorders and other abnormalities. Still, due to issues such as patient privacy, publicly available data is limited.
@@
As for methods to deal with the second problem in downstream tasks, transfer learning (Wang et al. 2023b) and meta-learning (Pati, Mewada, and Samanta 2023) have been proposed by reusing other downstream tasks from similar domains. They demonstrated the effectiveness of BCI domains, such as motor imagination. However, these approaches require a fine-tuned model trained on similar downstream tasks. Hence, they are hard to apply in cases where there is no sufficiently labeled data or fine-tuned model, such as in disease prediction.
@@
We focused on Parameter-Efficient Fine-Tuning (PEFT), which aims to fine-tune models with the competitive representation ability as a fine-tuned model with entire parameters with limited computational cost and datasets. The main idea of PEFT is to add a lightweight module as an ”adapter” with a pre-trained backbone model and fine-tune only the adapter while fixing the backbone model in each downstream task.
@@
Inspired by the concept of PEFT, we propose an EEG-GraphAdapter (EGA) to complement the understanding of spatial information in pre-trained models capturing temporal representations of EEG signals to overcome issues (1) and (2). EGA is a GNN-based module specially designed for downstream tasks by incorporating spatial representations into the input multivariate EEG signals to address the limitations of pre-trained models that only represent the temporal information of EEG. The BENDR model, pre-trained using stacked convolutional layers on large-scale EEG data, already captures the temporal characteristics of the EEG signals and remains unchanged during fine-tuning. This approach allows for efficient learning of spatial features without altering the pre-trained model, enabling effective adaptation to downstream tasks.
@@
EEG signals are typically recorded using electrodes placed at designated locations, and each electrode captures the time-series signals in a different area. By taking into account spatial information about the spatial representation of these electrodes, EEG data can reflect the temporal dynamics and spatial relationships of different brain regions. Many representation models handle data from individual EEG signals independently, without considering sensor location information or the spatial relationship between sensors. Spatial EEG information between sensors is essential for downstream tasks such as neurological disorder prediction and emotion recognition. For example, in autism spectrum disorder (ASD), abnormal functional connectivity between the frontal and parietal lobes has been observed (Coben, Mohammad-Rezazadeh, and Cannon 2014). Several GNN (Graph Neural Network) models that learn graph structures have been proposed as a model that encodes spatial relationships between EEG sensors for prediction tasks(Sun et al. 2024; Wagh and Varatharajah 2020; Demir et al. 2021).
@@
Integrating GNNs with time-series models such as Transformers makes it possible to capture the temporal dynamics of EEG signals and the spatial relationships between sensors (Wang et al. 2024; Yin et al. 2021). This fusion model addresses the limitations of traditional time-series models in EEG analysis.
The main contributions of this paper are as follows.
@@
  • We proposed EEG-GraphAdapter (EGA) with a GNN-based module to capture the spatial features between EEG sensors and enhance the pre-trained model BENDR representation learning capabilities with smaller computational costs and datasets.
@@
  • We conducted experiments on healthcare-related downstream tasks of MDD (Major Depressive Disorder) and Abnormality Detection. We demonstrated model and runtime performance improvements over the baseline BENDR model without the GraphAdapter.
Methodology
Model Architecture
Figure 1 shows the architecture diagram for pre-training the time-series model BENDR as a backbone and then adding the proposed method EGA to perform fine-tuning and prediction for the downstream task. The BENDR encodes EEG signals into a representation embedding for each sensor. It comprises six stacked 1D convolutional layers to extract distinctive features for EEG signals. This model is pre-trained using self-supervised learning techniques, such as random-masked sequence prediction, on multivariate EEG signals. During downstream tasks, such as the motor imaginary classification, parameters of the BENDR model are frozen. Linear Aggregator aggregates all BENDR encoded embeddings into a single representational embedding.
@@
Figure 1: Model Architectures in Pre-training and Downstream Tasks
EEG-GraphAdapter
@@
To incorporate spatial features into the input features for fine-tuning downstream tasks, we introduce EEG-GraphAdapter (EGA) and insert it before the BENDR model. The objective of the EGA is to predict a label $y\in{y_{1},y_{2},\ldots}$ from the multivariate EEG signals $X\in\{X_{1},X_{2},\ldots X_{n}\}=R^{L\times n}$, where $L$ and $n$ denotes the length of EEG signal and number of EEG sensors (channels), respectively. We define a fully connected weighted graph $G=(V,E)$ to represent the relationships between EEG sensors. $V$ denotes the set of EEG sensors ($|V|=n$), and $E=\{V\times V\}$ denotes the set of edges between sensors. We adopt the geodesic distance as the edge weight, using the same approach as EEG-GCNN(Wagh and Varatharajah 2020).
@@
Depending on the downstream task, many use EEG sequence samples with very short sequence lengths. While the BENDR and Transformer models can handle variable-length sequence data as is, the GNN model generally requires that the input feature Embedding have a fixed length. For this reason, raw EEG sequence data is preprocessed to the sequence length accepted by EGA using a linear transformation with a trainable linear layer.
@@
EGA comprises a two-layer GNN model representing spatial relationships between EEG sensors. The GNN processes the temporal raw EEG signal from each sensor $X$ and generates embeddings $X^{\prime}$ incorporating information from other sensors. In this framework, $GNN$ is represented by the EGA, and $A$ is the weighted adjacency matrix. We explored Graph Convolutional Network (GCN)(Kipf and Welling 2017), GraphSAGE(Hamilton, Ying, and Leskovec 2017), and Graph Attention Network (GAT)(Veličković et al. 2018) as GNN modules.
@@
The GCN(Kipf and Welling 2017) model aggregates information by averaging the feature embeddings $X_{j}$ of adjacent nodes (EEG sensors). In EGA, this approach allows for integrating EEG signal data across sensors, facilitating the learning of sensor relationships. Because the EEG graph is fully connected, embedding the EEG signal for each sensor aggregates the representation vectors of all other sensors except itself.
Besides, GraphSAGE(Hamilton, Ying, and Leskovec 2017) aggregates information from neighboring vertices (in this case, all sensors except the target sensor) by randomly sampling from the adjacent vertices.
GAT(Veličković et al. 2018) model adopts an ”attention mechanism” that assigns different weights (attention scores) to each neighboring node while aggregating their features. Unlike GCN and GraphSAGE, the GAT model learns individual weights for each node pair, automatically acquiring which sensor relationships are essential to identify labels in each downstream task.
BENDR
BENDR is a self-supervised model that captures EEG signals. Using stacked convolutional encoders, BENDR encodes the raw EEG signals into low-dimensional feature embeddings. The primary goal of BENDR is to enable various downstream tasks with limited computation resources and datasets by pre-training EEG signal data with a self-supervised learning approach. This approach allows for efficient learning and inference with limited data. While Transformer-based models are typically suited for handling time-series data, the computational cost of training on long EEG sequences is prohibitive. To address this, BENDR applies stacked 1D convolutional layers to the raw EEG signals, compressing the sequences into shorter embeddings before passing them into a Transformer model. BENDR extends the principles of wav2vec 2.0 (Baevski et al. 2020), which uses a Convolutional Neural Network (CNN) to reduce the dimension of raw audio data before applying a Transformer for pre-training. BENDR adapts this approach to the multi-channel nature of EEG data.
Although the BENDR model was designed to mitigate the computational cost associated with Transformers on long time-series sequences, when trying to fine-tune a combined model with the BENDR and a GNN in downstream tasks, the number of trainable parameters increases. As a result, additional computational costs and a large amount of data in each downstream task will be required for model convergence. Furthermore, there is a risk that the representational ability acquired during pre-training on large EEG datasets could be diminished during fine-tuning on specific downstream tasks, a phenomenon known as catastrophic forgetting.
BENDR encodes the original input EEG signals $X\in\mathbb{R}^{L\times n}$ into low-dimensional feature vectors $\tilde{X}\in\mathbb{R}^{d}$ ($d$ is the dimension of the encoded vector). During pre-training, the encoded feature vectors are masked, as shown on the left side of Figure 1, and the model employs self-supervised learning to reconstruct the masked features from the unmasked elements. The masked vectors pass through a Transformer encoder, which predicts the original unmasked vectors. The model parameters for both the BENDR and the Transformer encoder are updated using a self-supervised loss function based on the similarity between the predicted and original vectors.
Linear Aggregator and Classifier
Each downstream classification task outputs the prediction result from the representation vectors $\tilde{X}\in\mathbb{R}^{d}$, which the BENDR model encodes with Linear Aggregator and Classifier. First, the Linear Aggregator splits $\tilde{X}$ into four consecutive sub-vectors, and each sub-vector is averaged to produce a fixed-length representation vector $X_{out}\in\mathbb{R}^{4}$ (referred to as the Linear Aggregator). A linear layer with a softmax activation function is applied as a classifier to generate the final predictions.
The original BENDR paper explored an alternative approach where the Transformer encoder, also used during pre-training, was employed instead of the Linear Aggregator. However, the Linear Aggregator achieved better model performance than the Transformer encoder in most downstream tasks. Since the downstream tasks in this study focus on classification based on overall EEG waveform patterns, we consistently utilize the Linear Aggregator.
Experiments
To demonstrate that our proposed EGA performs classification downstream tasks with EEG data, we conduct the following experiments to answer the research questions:
@@
  • Which is the most effective among GNN models for the GraphAdapter?
Experimental Setup
@@
We adopted the BENDR model as the pre-trained model that forms the backbone of EGA. We pre-trained it from scratch using the latest Temple University EEG Corpus (TUEG) dataset (version 2.0.1, a total of 1.7TB, 69,652 samples) for our experiment. Our proposed EGA can be applied with other pre-trained models handling EEG signals if available.
The TUEG dataset includes EEG data recorded from various devices and patients. We applied preprocessing to make the data suitable for BENDR pre-training. First, we selected only EEG samples that meet the standard 10-20 sensor configuration, excluding samples that did not contain these channels. Additionally, we removed noise, such as artifacts from eye movements and the electric noise from devices. Specifically, we applied 50 Hz and band-pass filters in the 0.1 to 100 Hz range. Then, we standardized the sampling frequency to 256 Hz for all EEG samples.
@@
To evaluate the performance of EGA, we conducted model performance evaluations on two binary classification tasks (MDD, TUAB) related to healthcare as downstream tasks, as shown in Table 1. We used datasets (Mumtaz 2016; López et al. 2015) that are publicly available and have explicit labels for abnormal (positive) and healthy (negative) samples for downstream tasks.
@@
(Mumtaz 2016) This task predicts patients with Major Depressive Disorder (MDD; positive) and healthy control subjects (negative) based on EEG data.
@@
(López et al. 2015) It distinguishes whether the EEG sample is abnormal (positive) with sustained spikes or patterns such as Periodic Lateralized Epileptiform Discharges (PLEDs) or Generalized Periodic Epileptiform Discharges (GPEDs), or normal (negative) with the Posterior Dominant Rhythm (PDR) typically appears when a subject is relaxed with their eyes closed.
The MDD dataset (López et al. 2015) contains 19-channel EEG signals from patients with Major Depressive Disorder (MDD: positive) and healthy control (negative) subjects. For each subject, there are EEG samples measured during rest with eyes open (EO), eyes closed (EC), and a task in which the subject responds to letters displayed on the screen. In this experiment, we split the samples from EO and EC signals into 60-second segments and used them as sample data for evaluation. TUAB (López et al. 2015) involves predicting whether each EEG sample from The TUH Abnormal EEG Corpus is labeled normal (negative) or abnormal (positive). Given the imbalance in the dataset, we randomly selected 20 patients for each label and split each patient’s EEG sequence into 60-second segments for evaluation.
Name Task Length (s) Freq(Hz) #Subjects #EEG Samples %Positive Folds
@@
Table 1: Summary of Downstream Task Datasets
@@
Our experiments are conducted on an instance of our cloud computing platform named ”mdx”, with a single NVIDIA A100 GPU (40GB), 2-core Intel Xeon Platinum 8368 CPUs (38 cores, 2.4 GHz), and 512GiB DRAM.
We implement EGA as the extension of BENDR, which is implemented in  [^1] https://github.com/SPOClab-ca/BENDR with PyTorch backend. We also implement GNN models (GCN, GraphSAGE, and GAT) as the EGA modules in PyTorch Geometric. These GNN models have two hidden layers of size 64, and the number of attention heads was set to one for the GAT model. We used the Adam for the optimizer, with a fixed learning rate of 0.00001.
@@
Performance of EGA
To validate the effectiveness of the EGA, we compared the results of each downstream task against the baseline BENDR model. Figure 2 illustrates the architectural comparison of the models used for downstream tasks. The left side represents the baseline architecture, which consists of a trainable BENDR model followed by a Linear Aggregator. This architecture is identical to one of the models used in the original BENDR experiments for downstream tasks.
On the right side, the architecture shows the proposed method, where the parameters of the pre-trained BENDR model are fixed, and the EGA is introduced. In this configuration, the input includes time-series EEG data from 19 channels and a fully connected weighted EEG graph, where the distance between channels determines the weights. We adopt graph convolutional network (GCN), GraphSAGE, and Graph Attention Network (GAT) models as the EEG-GraphAdapter.
Figure 2: Comparison of Model Architectures in Downstream Tasks

### v2

# Graph Adapter for Parameter-Efficient Fine-Tuning of EEG Foundation Models

@@ Abstract
In diagnosing neurological disorders from electroencephalography (EEG) data, foundation models such as Transformers have been employed to capture temporal dynamics. Additionally, Graph Neural Networks (GNNs) are critical for representing the spatial relationships among EEG sensors. However, fine-tuning these large-scale models for both temporal and spatial features can be prohibitively large in computational cost, especially under the limited availability of labeled EEG datasets. We propose EEG-GraphAdapter (EGA), a parameter-efficient fine-tuning (PEFT) approach designed to address these challenges. EGA is integrated into a pre-trained temporal backbone model as a GNN-based module, freezing the backbone and allowing only the adapter to be fine-tuned. This enables the effective acquisition of EEG spatial representations, significantly reducing computational overhead and data requirements. Experimental evaluations on two healthcare-related downstream tasks—Major Depressive Disorder (MDD) and Abnormality Detection (TUAB)—show that EGA improves performance by up to 16.1% in F1-score compared with the backbone BENDR model, highlighting its potential for scalable and accurate EEG-based predictions.
Keywords:
Foundation Models, Graph Neural Networks, Parameter-efficient Fine-tuning
1 Introduction
@@
Electroencephalography (EEG) is a widely used method for measuring brain activity, capturing electrical signals from various brain regions through electrodes. While other imaging methods, such as functional MRI (fMRI), offer complementary insights into brain function, EEG uniquely provides direct and real-time access to brain dynamics. These signals reveal information related to functions such as consciousness, cognition, and motor activity, with applications spanning the healthcare field, including the diagnosis of mental disorders and the development of brain–computer interfaces (BCIs).
@@
As such, developing a foundation model for analyzing EEG data across diverse downstream tasks is essential in this domain. In recent years, with the development of large-scale neural network models such as LLMs, many foundation models for time series data have been proposed, including Transformer-based models. Methods for applying these foundation models to EEG signals have also been proposed bendr ; brainbert .
@@
However, there are several technical challenges in learning EEG representations: (1) Prohibitively large computational cost: As the number of parameters for fine-tuning in each downstream task increases, the computational resources for downstream tasks become more significant. (2) Insufficient training data: The available labeled EEG data sets are limited for many downstream tasks particularly in the healthcare domain due to the considerable effort required for data acquisition, measurement, and labeling by doctors. If fine-tuning is performed with a limited dataset, problems such as overfitting may arise, preventing the model from achieving sufficient accuracy in the downstream task. In particular, the demand for predictions is increasing in healthcare-related domains that predict neurological disorders and other abnormalities. However, due to issues such as patient privacy, publicly available data is limited.
@@
As for methods to deal with the second problem in downstream tasks, meta-learning meta-learning and transfer learning transfer-learning have been proposed to leverage knowledge from other similar domains and tasks. They demonstrated the effectiveness of BCI domains, such as motor imagination. However, these approaches require a fine-tuned model trained on similar downstream tasks. Hence, they are hard to apply in cases where there is no sufficiently labeled data or fine-tuned model, such as in disease prediction.
@@
We focus on Parameter-Efficient Fine-Tuning (PEFT), which aims to fine-tune models to achieve representation abilities comparable to fully fine-tuned models, while using fewer computational resources and less data. The main idea of PEFT is to add a lightweight module as an ”adapter” with a pre-trained backbone model and fine-tune only the adapter while fixing the backbone model in each downstream task.
@@
Inspired by the concept of PEFT, we propose EEG-GraphAdapter (EGA) to complement the understanding of spatial information in pre-trained models that capture the temporal representations of EEG signals, thereby addressing issues (1) and (2). EGA is a GNN-based module specially designed for downstream tasks by incorporating spatial representations into the input multivariate EEG signals to address the limitations of pre-trained models that only represent the temporal information of EEG.
The BENDR model, pre-trained using stacked convolutional layers on large-scale EEG data, already captures the temporal characteristics of the EEG signals and remains frozen during fine-tuning. This approach allows for efficient learning of spatial features without altering the pre-trained model, enabling effective adaptation to downstream tasks.
EEG signals are typically recorded using electrodes placed at designated locations, and each electrode captures the time-series signals in a different area. By taking into account the spatial arrangement of these electrodes, EEG data can reflect the temporal dynamics and spatial relationships of different brain regions.
@@
Many representation models handle data from individual EEG signals independently, without considering sensor location information or the spatial relationship between sensors. Spatial EEG information between sensors is essential for downstream tasks such as neurological disorder prediction and emotion recognition. For example, in autism spectrum disorder (ASD), abnormal functional connectivity between the frontal and parietal lobes has been observed asd1 . Several GNN (Graph Neural Network) models that learn graph structures have been proposed as a model that encodes spatial relationships between EEG sensors for prediction taskseeg-gcnn ; eeg-gnn ; mggcn .
@@
Integrating GNNs with time-series models such as Transformers makes it possible to capture the temporal dynamics of EEG signals and the spatial relationships between sensors erdl ; spatio-temporal-eeg . This fusion model addresses the limitations of traditional time-series models in EEG analysis. The main contributions of this paper are as follows.
@@
  • We propose EEG-GraphAdapter (EGA), a GNN-based module that captures the spatial features between EEG sensors and enhances the representation learning capabilities of the pre-trained BENDR model, while requiring fewer computational resources and less data.
@@
  • We conducted experiments on two healthcare-related downstream tasks, MDD (Major Depressive Disorder) and Abnormality Detection, demonstrating improvements in both model performance and runtime compared to the baseline BENDR model without the GraphAdapter.
2 Methodology
2.1 Model Architecture
Figure 1 illustrates the architecture for pre-training the time-series BENDR model as the backbone (Left) and for the downstream task using the baseline BENDR model (Center) and the proposed method EGA (Right).
@@ 2.1 Model Architecture
Figure 1: Model Architectures in Pre-training, and Downstream Tasks with Fully Fine-tuned BENDR and Proposed EGA with BENDR Frozen
BENDR Pre-training The left subfigure illustrates BENDR’s self-supervised pre-training process. The parameters of the BENDR and Transformer models (Transformer Encoder) are updated by a random-masked sequence prediction task, respectively. The encoded feature vectors are masked, and the model employs self-supervised learning to reconstruct the masked features from the unmasked elements. The masked vectors pass through a Transformer Encoder, which predicts the original unmasked vectors. The model parameters for both the BENDR and the Transformer Encoder are updated using a self-supervised loss function based on the similarity between the predicted and original vectors.
BENDR Fine-tuning (Baseline) The center subfigure shows the baseline approach for downstream fine-tuning, which is consistent with the original BENDR experiments. We replace the Transformer Encoder with the simpler Linear Aggregator, which better suits classification tasks. Here, the pre-trained BENDR model is fully fine-tuned on the classification task, and its output embeddings are combined into a single feature vector via the Linear Aggregator before the final prediction.
EGA Fine-tuning (Proposed) The right subfigure shows our proposed method, EEG-GraphAdapter (EGA). In contrast to the baseline approach, the BENDR model is frozen to preserve the representation learned during pre-training. Instead, the EGA module is introduced before the BENDR to incorporate spatial relationships between EEG sensors, improving model performance without fine-tuning the BENDR model.
The following sections describe the BENDR, Linear Aggregator, and EGA used in the proposed method.
BENDR
BENDR is a self-supervised model designed to capture the unique characteristics of EEG signals. It comprises six stacked 1D convolutional layers to extract meaningful features from raw EEG signals. It encodes the original EEG signals $X\in\mathbb{R}^{L\times n}$ into low-dimensional representation feature vectors $\tilde{X}\in\mathbb{R}^{d}$ ($d$ is the dimension of the encoded vector).
BENDR is pre-trained on multivariate EEG signals using self-supervised learning techniques, such as random-masked sequence prediction, to learn robust data representations. The primary objective of BENDR is to facilitate various downstream tasks, even in scenarios with limited computational resources and small datasets. By leveraging self-supervised pre-training, BENDR enables efficient learning and inference, making it well-suited for tasks with constrained data availability.
While Transformer-based models are typically suited for handling time-series data, the computational cost of training on long EEG sequences can be extremely high. To address this, BENDR applies stacked 1D convolutional layers to the raw EEG signals, compressing the sequences into shorter embeddings before passing them into a Transformer model. BENDR extends the principles of wav2vec 2.0 wav2vec2 , which uses a Convolutional Neural Network (CNN) to reduce the dimension of raw audio data before applying a Transformer for pre-training. BENDR adapts this approach to the multi-channel nature of EEG data.
Although the BENDR model was designed to mitigate the computational cost associated with Transformers on long time-series sequences, when trying to fine-tune a combined model with the BENDR and a GNN in downstream tasks, the number of trainable parameters increases. As a result, additional computational costs and a large amount of data in each downstream task will be required for model convergence. Furthermore, there is a risk that the representational ability acquired during pre-training on large EEG datasets could be diminished during fine-tuning for specific downstream tasks, a phenomenon known as catastrophic forgetting.
Linear Aggregator
Linear Aggregator aggregates all embeddings encoded by BENDR into a single representation in downstream tasks. Despite the Transformer Encoder’s integral role in pre-training, experimental results in the original BENDR paperbendr show that a simpler Linear Aggregator often yields better performance than the Transformer Encoder when fine-tuning on downstream classification tasks. Consequently, while the Transformer Encoder is necessary during pre-training to learn high-level contextual representations via masked sequence prediction, it is not necessarily optimal to include it in the final classification pipeline. Since the downstream tasks in this study focus on classification based on overall EEG waveform patterns, we consistently utilize the Linear Aggregator.
Each downstream classification task outputs predictions from the representation vectors $\tilde{X}\in\mathbb{R}^{d}$, which are produced by the BENDR model and then passed through the Linear Aggregator and Classifier. First, the Linear Aggregator splits $\tilde{X}$ into four consecutive sub-vectors, and each sub-vector is averaged to produce a fixed-length representation vector $X_{out}\in\mathbb{R}^{4}$ (referred to as the Linear Aggregator). A linear layer with a softmax activation function is applied as a classifier to generate the final predictions.
EEG-GraphAdapter (EGA)
@@
To incorporate spatial features into the input features for fine-tuning downstream tasks, we introduce EEG-GraphAdapter (EGA) and place it before the BENDR model. The objective of the EGA is to predict a label $y\in{y_{1},y_{2},\ldots}$ from the multivariate EEG signals $X\in\{X_{1},X_{2},\ldots X_{n}\}=R^{L\times n}$, where $L$ and $n$ denote the length of the EEG signal and number of EEG sensors (channels), respectively. We define a fully connected weighted graph $G=(V,E)$ to represent the relationships between EEG sensors. $V$ denotes the set of EEG sensors ($|V|=n$), and $E\subseteq\{V\times V\}$ denotes the set of edges between sensors. We adopt the geodesic distance as the edge weight, using the same approach as EEG-GCNNeeg-gcnn .
@@
Depending on the downstream task, EEG sequence samples often have very short sequence lengths. While the BENDR and Transformer models can handle variable-length sequence data as is, the GNN model generally requires the input feature embedding to have a fixed length. For this reason, the raw EEG sequence data is preprocessed to the sequence length required by EGA using a trainable linear layer.
@@
EGA comprises a two-layer GNN model representing spatial relationships between EEG sensors. The GNN processes the temporal EEG signals from each sensor $X$ and generates embeddings $X^{\prime}$ that incorporate information from other sensors. In this framework, $GNN$ is represented by the EGA, and $A$ is the weighted adjacency matrix. We explored Graph Convolutional Network (GCN)gcn , GraphSAGEgraphsage , and Graph Attention Network (GAT)gat as GNN modules.
@@
The GCNgcn model aggregates information by averaging the feature embeddings $X_{j}$ of adjacent nodes (EEG sensors). In EGA, this approach allows for integrating EEG signal data across sensors, facilitating the learning of sensor relationships. Because the EEG graph is fully connected, embedding the EEG signal for each sensor aggregates the representation vectors of all other sensors except the sensor itself.
Besides, GraphSAGEgraphsage aggregates information from neighboring vertices (in this case, all sensors except the target sensor) by randomly sampling from the adjacent vertices. GATgat model adopts an ”attention mechanism” that assigns different weights (attention scores) to each neighboring node while aggregating their features. Unlike GCN and GraphSAGE, the GAT model learns individual weights for each node pair, automatically determining which sensor relationships are crucial in predicting labels for each downstream task.
3 Experiments
To demonstrate that our proposed EGA can handle classification on downstream tasks using EEG data, we conduct the following experiments to answer the research questions:
@@ 3 Experiments
  • Which GNN model is the most effective for the GraphAdapter?
3.1 Experimental Setup
@@
We adopted the BENDR model as the pre-trained model that forms the backbone of EGA. We pre-trained it from scratch using the latest Temple University EEG Corpus (TUEG) dataset (version 2.0.1, a total of 1.7TB, 69,652 samples) for our experiment. Our proposed EGA can be applied with other pre-trained models handling EEG signals where available.
The TUEG dataset includes EEG data recorded from various devices and patients. We applied preprocessing to make the data suitable for BENDR pre-training. First, we selected only EEG samples that meet the standard 10-20 sensor configuration, excluding samples that did not contain these channels. Additionally, we removed noise, such as artifacts from eye movements and electrical noise from devices. Specifically, we applied 50 Hz and band-pass filters in the 0.1 to 100 Hz range. Then, we standardized the sampling frequency to 256 Hz for all EEG samples.
@@
To evaluate the performance of EGA, we conducted binary classification experiments on two healthcare-related downstream tasks (MDD and TUAB), as shown in Table 1. We used datasets mdd-data ; tuab-data that are publicly available and have explicit labels for abnormal (positive) and healthy (negative) samples for downstream tasks.
@@
mdd-data This task predicts patients with Major Depressive Disorder (MDD; positive) and healthy control subjects (negative) based on EEG data.
@@
tuab-data It distinguishes whether the EEG sample is abnormal (positive) with sustained spikes or patterns such as Periodic Lateralized Epileptiform Discharges (PLEDs) or Generalized Periodic Epileptiform Discharges (GPEDs), or normal (negative) with the Posterior Dominant Rhythm (PDR) typically appears when a subject is relaxed with their eyes closed.
The MDD dataset mdd-data contains 19-channel EEG signals from patients with Major Depressive Disorder (MDD: positive) and healthy control (negative) subjects. For each subject, there are EEG samples measured during rest with eyes open (EO), eyes closed (EC), and a task in which the subject responds to letters displayed on the screen. In this experiment, we split the samples from EO and EC signals into 60-second segments and used them as sample data for evaluation. TUAB tuab-data involves predicting whether each EEG sample from The TUH Abnormal EEG Corpus is labeled normal (negative) or abnormal (positive). Given the imbalance in this dataset, we randomly selected 20 patients for each label and split each patient’s EEG sequence into 60-second segments for evaluation.
Name Task Length (s) Freq(Hz) #Subjects #Samples %Positive Folds
@@
Table 1: Datasets for Downstream Tasks
@@
Our experiments were conducted on an instance of our academic cloud platform named ”mdx”mdx , equipped with a single NVIDIA A100 GPU (40GB), two Intel Xeon Platinum 8368 CPUs (each with 38 cores at 2.4 GHz), and 512 GiB of DRAM.
We implement EGA as the extension of BENDR, which is implemented in  [^1] https://github.com/SPOClab-ca/BENDR with PyTorch backend. We also implement GNN models (GCN, GraphSAGE, and GAT) as EGA modules in PyTorch Geometric. These GNN models have two hidden layers of size 64, and the number of attention heads was set to one for the GAT model. We used Adam for the optimizer, with a fixed learning rate of 0.00001.
@@
3.2 Performance of EGA
To validate the effectiveness of the EGA, we compared the results of each downstream task against the baseline BENDR model. The architectures of the baseline and the proposed EGA are illustrated in Figure 1. In this configuration, the input includes time-series EEG data from 19 channels and a fully connected weighted EEG graph, where the distance between channels determines the weights. We adopt graph convolutional network (GCN), GraphSAGE, and Graph Attention Network (GAT) models as the EEG-GraphAdapter.